OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (-2, -1) and has a slope of 5.

OpenStudy (anonymous):

y - 1 = 5(x - 2) y + 1 = 5(x + 2) y + 2 = 5(x + 1) y - 2 = 5(x - 1)

OpenStudy (precal):